# Asymmetric Errors

### v1

@@ 3.2 Combination of results using likelihood
Likelihood-based errors are greatly used in the combination of results. The combination of errors will be dealt with in the Section 3.4..
If one has several results measuring the same quantity then the total likelihood is the product of the individual likelihoods, and the log likelihood is just their sum. From that one can read off the overall $\hat{a}$ and the $\Delta\ln L=-{1\over 2}$ errors. If we do not know the full forms for the likelihoods, but are only given the peak values and errors, then we can use the same approach, but modelling the likelihoods by one of the methods above. In general a numerical maximisation can be used to give a solution. For the two Bartlett models this can be done efficiently: the likelihood is
@@
The equivalent for the linear Variance model is
@@
Solutions can be found by iteration. In all cases one can then find the errors on the final result by numerical interpolation, to discover where the log likelihood falls by ${1\over 2}$ from its peak value.
An illustration is shown in Figure 12. The 3 values ${1.9}^{+0.7}_{-0.5}$, ${2.4}^{+0.6}_{-0.8}$ and ${3.1}^{+0.5}_{-0.4}$ are combined (using the linear variance model) to give ${2.754}^{+0.286}_{-0.263}$. For the linear sigma model the result is ${2.758}^{+0.293}_{-0.272}$ and the plot is identical to the eye. Indeed other models give very similar results, as shown in Table LABEL:tab:combineresults. The likelihood plots all look very similar to Figure 12, apart from that using the cubic model.
@@ 3.3 Goodness of fit with likelihood
According to Wilks’ theorem, under the null hypothesis the improvement in likelihood due to the introduction of $n$ extra parameters is such that $-2\Delta\ln L$ is distributed according to a $\chi^{2}$ distribution. On that basis, the change in the log likelihood when using only 1 free parameter (the combined result) rather then $n$ fitted values (the individual results) gives a measure of goodness of fit such that $-2\Delta\ln L$ can is $\chi^{2}$ with $n-1$ degrees of freedom.
Wilks’ theorem holds only in the limit of large numbers, and provided the first model contains the second. The first point depends clearly on the number of results being combined. The second is debatable. So this needs to be investigated empirically.
@@ 3.4 Combination of errors with likelihood
For the linear sigma model, $w_{i}={\left(\sigma_{i}+\sigma^{\prime}_{i}a_{i}\right)^{3}\over\sigma_{i}}$, for the linear variance model $w_{i}={\left(V_{i}+a_{i}V^{\prime}_{i}\right)^{2}\over 2V_{i}+V^{\prime}_{i}a_{i}}$.
@@ 4.1.1 A signal with several backgrounds: combination of errors using likelihoods
The same total background measurement of 9 might be obtained in various ways, such as 3+6 or 3+3+3 or even 9 separate measurements of 1 event. The same combination procedure also gives ${}^{+3.333}_{-2.668}$ for the 3+6 case, and changes only slightly to ${}^{+3.323}_{-2.659}$ for 3+3+3 and ${}^{+3.269}_{-2.610}$ for 1+1+1+1+1+1+1+1+1. Outcomes for other models and for other inputs are shown in Table LABEL:tab:coelike, and show remarkable agreement with the true values for the errors.
4.1.2 A lifetime measurement
@@
Suppose an experiment measures a lifetime $\tau$ from 3 decays. The values happen to be 1.241, 0.592, and 0.988, in some time units. (These numbers were generated from an exponential distribution, ensuring that this example is realistic.) Maximising $\ln L$ (which is $\ln{\rm exp}(-t/\tau)/\tau$) and using $\Delta\ln L=-{1\over 2}$ gives a result $\tau={0.940}^{+0.841}_{-0.385}$.
@@
The experiment then measures 3 more lifetimes, which happen to be 0.834, 2.964, and 0.176, which combined on their own give ${1.325}^{+1.184}_{-0.542}$. If we combine all 6 values we get the best result ${1.1325}^{+0.6225}_{-0.3598}$. These are all ‘correct’ values in the sense that they use the fact that the likelihood is exponential.
@@
Now suppose we take the two partial results separately (i.e. just the value and $\pm$ errors, as quoted above) and combine them using the linear-sigma model. That gives ${1.1323}^{+0.6213}_{-0.3604}$. The linear-variance model gives ${1.1318}^{+0.6249}_{-0.3577}$. Both these agree well with the ‘correct’ value, both in the central value and the quoted errors. Both models (which know nothing about the fact that this was a lifetime measurement with an exponential likelihood) give a result very close to the full-information ‘correct’ one.
@@
These results are shown in Table LABEL:tab:life together with those from some other models. The Edgeworth model cannot cope with asymmetries this large, and the cubic does not find a positive error due to the turnover in the curve. It can be seen that the logarithmic and PDG models also do well, the quartic is less good, and the Poisson and Azzalini models perform relatively poorly. The final row, labelled “wrong”, takes the mean of the two results and combines the positive and negative errors separately in quadrature, and it can be seen that its error estimates are seriously discrepant.
4.1.3 Combining Poisson counts
@@
Suppose a counting experiment sees 5 events in an hour. The result is quoted (using $\Delta\ln L=-{1\over 2}$ errors, even though this is a case where the full Neyman errors could be given) as ${5}^{+2.581}_{-1.916}$. This continues for another hour and, as it happens, 5 events are again seen. The total gives a result ${10}^{+3.504}_{-2.838}$ and with the knowledge we have of the way the experiment has been done, we can estimate the number of events per hour by dividing this by 2 to get ${5}^{+1.752}_{-1.419}$.
@@
Table 6: Combining results from two samples from the same Poisson distribution. The ideal result is ${5.000}^{+1.752}_{-1.419}$

### v2

@@ 3.2 Combination of results using likelihood
Likelihood-based errors are greatly used in the combination of results. The combination of errors will be dealt with in the Section 3.4.
If one has several results measuring the same quantity then the total likelihood is the product of the individual likelihoods, and the log likelihood is just the sum of the individual log likelihoodsFrom that one can read off the overall $\hat{a}$ and the $\Delta\ln L=-{1\over 2}$ errors. If we do not know the full forms for the likelihoods, but are only given the peak values and errors, then we can use the same approach, but modelling the likelihoods by one of the methods above. In general a numerical maximisation can be used to give a solution. For the two Bartlett models this can be done efficiently: the likelihood is
@@
The equivalent for the linear v‘ariance model is
@@
Solutions can be found by iteration. In all cases one can then find the errors on the final result by numerical root finding, to discover where the log likelihood falls by ${1\over 2}$ from its peak value.
An illustration is shown in Figure 13. The 3 values ${1.9}^{+0.7}_{-0.5}$, ${2.4}^{+0.6}_{-0.8}$ and ${3.1}^{+0.5}_{-0.4}$ are combined (using the linear variance model) to give ${2.754}^{+0.286}_{-0.263}$. For the linear sigma model the result is ${2.758}^{+0.293}_{-0.272}$ and the plot is identical to the eye. Indeed other models give very similar results, as shown in Table 7. The likelihood plots all look very similar to Figure 13, apart from that using the cubic model.
Model Result Model Result
Cubic fails Truncated cubic ${2.792}^{+0.342}_{-0.306}$
Broken parabola ${2.703}^{+0.301}_{-0.301}$ Symmetrized parabola ${2.666}^{+0.321}_{-0.321}$
Constrained quartic ${2.765}^{+0.303}_{-0.285}$ Molded quartic ${2.721}^{+0.246}_{-0.240}$
Matched quintic ${2.728}^{+0.290}_{-0.300}$ Interpolated 7th degree ${2.702}^{+0.301}_{-0.296}$
Simple double quartic ${2.720}^{+0.289}_{-0.299}$ Molded double quartic ${2.718}^{+0.286}_{-0.298}$
Simple double quintic ${2.702}^{+0.301}_{-0.298}$ Molded double quintic ${2.701}^{+0.302}_{-0.297}$
Conservative spline(0.05) ${2.720}^{+0.291}_{-0.292}$ Conservative spline(0.1) ${2.732}^{+0.285}_{-0.287}$
Conservative spline(0.15) ${2.738}^{+0.285}_{-0.281}$ Conservative spline(0.20) ${2.742}^{+0.287}_{-0.281}$
Conservative spline(max) ${2.742}^{+0.287}_{-0.282}$ Log logistic beta ${2.763}^{+0.336}_{-0.366}$
Logarithmic ${2.755}^{+0.288}_{-0.266}$ Generalised Poisson ${2.753}^{+0.283}_{-0.258}$
Variable sigma ${2.758}^{+0.293}_{-0.272}$ Variable variance ${2.754}^{+0.286}_{-0.263}$
Variable log sigma ${2.743}^{+0.289}_{-0.291}$ Molded cubic log sigma ${2.702}^{+0.301}_{-0.297}$
Quintic log sigma ${2.702}^{+0.301}_{-0.296}$ PDG ${2.726}^{+0.273}_{-0.309}$
Skew normal ${2.749}^{+0.293}_{-0.285}$
Railway Gaussian ${2.753}^{+0.277}_{-0.250}$ Double cubic Gaussian ${2.731}^{+0.294}_{-0.316}$
Symmetric beta Gaussian(1, 10) ${2.745}^{+0.282}_{-0.314}$ Symmetric beta Gaussian(1, 30) ${2.751}^{+0.281}_{-0.260}$
Symmetric beta Gaussian(4, 10) ${2.701}^{+0.309}_{-0.309}$ Symmetric beta Gaussian(4, 30) ${2.748}^{+0.286}_{-0.278}$
QVW Gaussian ${2.741}^{+0.290}_{-0.297}$ Johnson system
Log normal ${2.755}^{+0.288}_{-0.266}$
Table 7: Results of combining the same 3 results according to different models.
@@
It is worth pointing out how the combined error depends on the values being combined, as well as their quoted errors. If one of the values in the combination is changed this affects not only the combined central value but also the quoted error. In the symmetric case the quoted errors are independent of the values.
@@ 3.3 Goodness of fit with likelihood
[…]
Wilks’ theorem holds only in the limit of large $N$and provided the first model contains the second. The first point depends clearly on the number and precision of theresults being combined. . So this needs to be investigated empirically.
@@ 3.4 Combination of errors with likelihood
For the linear sigma model, $w_{i}={\left(\sigma_{i}+\sigma^{\prime}_{i}a_{i}\right)^{3}\over\sigma_{i}}$, for the linear variance model $w_{i}={\left(V_{i}+a_{i}V^{\prime}_{i}\right)^{2}\over 2V_{i}+V^{\prime}_{i}a_{i}}$, and similarly for other models. If the cancellation cannot be done algebraically and numerical evaluation gives 0/0 at the peak, one can use $w=-\left({\partial^{2}L\over\partial a^{2}}\right)^{-1}$ for small $a$.
@@
Inputs Broken parabola Truncated cubic Constrained quartic Molded quartic Matched quintic Interpolated 7th degree
4 5 ${9}^{+3.488}_{-2.549}$ ${9}^{+3.218}_{-2.620}$ ${9}^{+3.272}_{-2.635}$ ${9}^{+3.439}_{-2.678}$ ${9}^{+3.283}_{-2.590}$ ${9}^{+3.425}_{-2.558}$
3 6 ${9}^{+3.483}_{-2.556}$ ${9}^{+3.224}_{-2.620}$ ${9}^{+3.274}_{-2.635}$ ${9}^{+3.433}_{-2.676}$ ${9}^{+3.275}_{-2.594}$ ${9}^{+3.402}_{-2.567}$
2 7 ${9}^{+3.471}_{-2.572}$ ${9}^{+3.238}_{-2.622}$ ${9}^{+3.278}_{-2.636}$ ${9}^{+3.417}_{-2.670}$ ${9}^{+3.264}_{-2.602}$ ${9}^{+3.345}_{-2.583}$
1 8 ${9}^{+3.450}_{-2.601}$ ${9}^{+3.264}_{-2.628}$ ${9}^{+3.290}_{-2.640}$ ${9}^{+3.368}_{-2.659}$ ${9}^{+3.269}_{-2.617}$ ${9}^{+3.279}_{-2.608}$
3 3 3 ${9}^{+3.603}_{-2.453}$ ${9}^{+3.105}_{-2.566}$ ${9}^{+3.206}_{-2.594}$ ${9}^{+3.500}_{-2.669}$ ${9}^{+3.161}_{-2.523}$ ${9}^{+3.356}_{-2.477}$
1 1 1 1 1 1 1 1 1 ${9}^{+4.073}_{-2.095}$ ${9}^{+2.605}_{-2.284}$ ${9}^{+2.894}_{-2.386}$ ${9}^{+3.389}_{-2.508}$ ${9}^{+2.552}_{-2.217}$ ${9}^{+2.576}_{-2.157}$
Inputs Double quartic Molded double quartic Double quintic Molded double quintic Conservative spline (0.05) Conservative spline (0.10)
4 5 ${9}^{+3.383}_{-2.600}$ ${9}^{+3.435}_{-2.617}$ ${9}^{+3.471}_{-2.558}$ ${9}^{+3.479}_{-2.561}$ ${9}^{+3.426}_{-2.592}$ ${9}^{+3.379}_{-2.618}$
3 6 ${9}^{+3.376}_{-2.606}$ ${9}^{+3.432}_{-2.623}$ ${9}^{+3.461}_{-2.568}$ ${9}^{+3.473}_{-2.573}$ ${9}^{+3.423}_{-2.596}$ ${9}^{+3.377}_{-2.620}$
2 7 ${9}^{+3.362}_{-2.617}$ ${9}^{+3.428}_{-2.635}$ ${9}^{+3.438}_{-2.588}$ ${9}^{+3.460}_{-2.595}$ ${9}^{+3.416}_{-2.606}$ ${9}^{+3.374}_{-2.626}$
1 8 ${9}^{+3.342}_{-2.636}$ ${9}^{+3.421}_{-2.651}$ ${9}^{+3.393}_{-2.619}$ ${9}^{+3.439}_{-2.627}$ ${9}^{+3.403}_{-2.624}$ ${9}^{+3.367}_{-2.636}$
3 3 3 ${9}^{+3.378}_{-2.552}$ ${9}^{+3.506}_{-2.589}$ ${9}^{+3.536}_{-2.481}$ ${9}^{+3.575}_{-2.492}$ ${9}^{+3.489}_{-2.526}$ ${9}^{+3.405}_{-2.569}$
1 1 1 1 1 1 1 1 1 ${9}^{+3.250}_{-2.380}$ ${9}^{+3.888}_{-2.509}$ ${9}^{+3.517}_{-2.245}$ ${9}^{+3.963}_{-2.306}$ ${9}^{+3.727}_{-2.263}$ ${9}^{+3.482}_{-2.355}$
Inputs Conservative spline (0.15) Conservative spline (0.20) Conservative spline (max) Log logistic-beta Logarithmic Generalized Poisson
4 5 ${9}^{+3.342}_{-2.628}$ ${9}^{+3.316}_{-2.627}$ ${9}^{+3.315}_{-2.627}$ ${9}^{+3.177}_{-2.545}$ ${9}^{+3.325}_{-2.663}$ ${9}^{+3.342}_{-2.676}$
3 6 ${9}^{+3.342}_{-2.629}$ ${9}^{+3.319}_{-2.629}$ ${9}^{+3.313}_{-2.628}$ ${9}^{+3.192}_{-2.555}$ ${9}^{+3.325}_{-2.663}$ ${9}^{+3.342}_{-2.676}$
2 7 ${9}^{+3.342}_{-2.632}$ ${9}^{+3.322}_{-2.632}$ ${9}^{+3.309}_{-2.631}$ ${9}^{+3.221}_{-2.576}$ ${9}^{+3.325}_{-2.663}$ ${9}^{+3.345}_{-2.583}$
1 8 ${9}^{+3.341}_{-2.640}$ ${9}^{+3.325}_{-2.640}$ ${9}^{+3.306}_{-2.638}$ ${9}^{+3.264}_{-2.609}$ ${9}^{+3.326}_{-2.663}$ ${9}^{+3.342}_{-2.676}$
3 3 3 ${9}^{+3.341}_{-2.585}$ ${9}^{+3.292}_{-2.584}$ ${9}^{+3.272}_{-2.582}$ ${9}^{+3.071}_{-2.459}$ ${9}^{+3.308}_{-2.649}$ ${9}^{+3.342}_{-2.676}$
1 1 1 1 1 1 1 1 1 ${9}^{+3.310}_{-2.383}$ ${9}^{+3.187}_{-2.386}$ ${9}^{+3.025}_{-2.370}$ ${9}^{+2.731}_{-2.178}$ ${9}^{+3.211}_{-2.572}$ ${9}^{+3.342}_{-2.676}$
Inputs Linear sigma Linear variance Linear log sigma Molded cubic log sigma Quintic log sigma PDG method
4 5 ${9}^{+3.310}_{-2.653}$ ${9}^{+3.333}_{-2.668}$ ${9}^{+3.322}_{-2.624}$ ${9}^{+3.475}_{-2.561}$ ${9}^{+3.443}_{-2.562}$ ${9}^{+3.310}_{-2.653}$
3 6 ${9}^{+3.310}_{-2.653}$ ${9}^{+3.333}_{-2.668}$ ${9}^{+3.318}_{-2.626}$ ${9}^{+3.467}_{-2.572}$ ${9}^{+3.427}_{-2.571}$ ${9}^{+3.311}_{-2.653}$
2 7 ${9}^{+3.311}_{-2.653}$ ${9}^{+3.333}_{-2.668}$ ${9}^{+3.311}_{-2.630}$ ${9}^{+3.449}_{-2.595}$ ${9}^{+3.390}_{-2.588}$ ${9}^{+3.311}_{-2.653}$
1 8 ${9}^{+3.313}_{-2.654}$ ${9}^{+3.333}_{-2.668}$ ${9}^{+3.304}_{-2.638}$ ${9}^{+3.416}_{-2.628}$ ${9}^{+3.315}_{-2.613}$ ${9}^{+3.313}_{-2.654}$
3 3 3 ${9}^{+3.278}_{-2.630}$ ${9}^{+3.323}_{-2.659}$ ${9}^{+3.278}_{-2.582}$ ${9}^{+3.555}_{-2.491}$ ${9}^{+3.437}_{-2.487}$ ${9}^{+3.278}_{-2.630}$
1 1 1 1 1 1 1 1 1 ${9}^{+3.098}_{-2.500}$ ${9}^{+3.269}_{-2.610}$ ${9}^{+2.971}_{-2.381}$ ${9}^{+3.735}_{-2.320}$ ${9}^{+2.910}_{-2.202}$ ${9}^{+3.098}_{-2.500}$
Inputs Skew normal Railway Gaussian Double cubic Gaussian Symmetric beta Gaussian(1,10) Symmetric beta Gaussian(1,30) Symmetric beta Gaussian(4,10)
4 5 ${9}^{+3.302}_{-2.628}$ ${9}^{+3.355}_{-2.705}$ ${9}^{+3.312}_{-2.579}$ ${9}^{+3.330}_{-2.626}$ ${9}^{+3.350}_{-2.678}$ ${9}^{+3.347}_{-2.518}$
3 6 ${9}^{+3.304}_{-2.630}$ ${9}^{+3.356}_{-2.707}$ ${9}^{+3.291}_{-2.584}$ ${9}^{+3.328}_{-2.627}$ ${9}^{+3.350}_{-2.679}$ ${9}^{+3.319}_{-2.528}$
2 7 ${9}^{+3.308}_{-2.635}$ ${9}^{+3.359}_{-2.711}$ ${9}^{+3.254}_{-2.597}$ ${9}^{+3.322}_{-2.630}$ ${9}^{+3.350}_{-2.680}$ ${9}^{+3.261}_{-2.548}$
1 8 ${9}^{+3.318}_{-2.645}$ ${9}^{+3.366}_{-2.720}$ ${9}^{+3.310}_{-2.634}$ ${9}^{+3.351}_{-2.683}$
3 3 3 ${9}^{+3.271}_{-2.591}$ ${9}^{+3.373}_{-2.740}$ ${9}^{+3.174}_{-2.509}$ ${9}^{+3.303}_{-2.587}$ ${9}^{+3.357}_{-2.683}$ ${9}^{+3.195}_{-2.399}$
1 1 1 1 1 1 1 1 1 ${9}^{+3.166}_{-2.444}$ ${9}^{+3.543}_{-3.032}$ ${9}^{+3.039}_{-2.350}$ ${9}^{+3.409}_{-2.740}$
Inputs Symmetric beta Gaussian(4,30) QVW Gaussian Johnson system Log normal
4 5 ${9}^{+3.341}_{-2.652}$ ${9}^{+3.330}_{-2.620}$ ${9}^{+3.325}_{-2.663}$
3 6 ${9}^{+3.339}_{-2.653}$ ${9}^{+3.326}_{-2.623}$ ${9}^{+3.325}_{-2.663}$
2 7 ${9}^{+3.336}_{-2.655}$ ${9}^{+3.318}_{-2.628}$ ${9}^{+3.325}_{-2.663}$
1 8 ${9}^{+3.332}_{-2.659}$ ${9}^{+3.308}_{-2.637}$ ${9}^{+3.326}_{-2.663}$
3 3 3 ${9}^{+3.330}_{-2.634}$ ${9}^{+3.293}_{-2.578}$ ${9}^{+3.308}_{-2.649}$
1 1 1 1 1 1 1 1 1 ${9}^{+3.238}_{-2.546}$ ${9}^{+3.013}_{-2.374}$ ${9}^{+3.211}_{-2.572}$
Table 8: Examples of combination of errors using likelihoods. The ‘correct’ answer is ${9}^{+3.342}_{-2.676}$. Where models cannot accommodate the asymmetry (the $N=1$ Poisson is strongly asymmetric and breaks several models, and the Johnson System cannot handle even moderate asymmetries) no result can be given.
@@ 4.1.1 A signal with several backgrounds: combination of errors using likelihoods
The same total background measurement of 9 might be obtained in various ways, such as 3+6 or 3+3+3 or even 9 separate measurements of 1 event. The same combination procedure also gives ${}^{+3.333}_{-2.668}$ for the 3+6 case, and changes only slightly to ${}^{+3.323}_{-2.659}$ for 3+3+3 and ${}^{+3.269}_{-2.610}$ for 1+1+1+1+1+1+1+1+1. Outcomes for other models and for other inputs are shown in Table 8, and show remarkable agreement with the true values for the errors.
4.1.2 A lifetime measurement: combination of results using likelihoods
@@
Suppose an experiment measures a lifetime $\tau$ from 3 decays. The values happen to be 1.241, 0.592, and 0.988, in some time units. (These numbers were generated from an exponential distribution, ensuring that this example is realistic.) Maximising $\ln L$, which is $\ln{\rm exp}(-t/\tau)/\tau$, and using $\Delta\ln L=-{1\over 2}$ gives a result $\tau={0.940}^{+0.841}_{-0.385}$.
@@
The experiment then measures 3 more lifetimes, which happen to be 0.834, 2.964, and 0.176, which combined on their own give ${1.325}^{+1.184}_{-0.542}$.
If we combine all 6 values we get the combinedresult ${1.1325}^{+0.6225}_{-0.3598}$. The likelihood used is the product of 6 separate exponential likelihoods, so we are using the full information..
@@
Now suppose we take the two partial results separately (i.e. just the value and $\pm$ errors, as quoted above) and combine them using the linear-sigma model. That gives ${1.1323}^{+0.6213}_{-0.3604}$. The linear-variance model gives ${1.1318}^{+0.6249}_{-0.3577}$. Both these agree well with the full informationvalue, both in the central value and the quoted errors. Both models (which know nothing about the fact that this was a lifetime measurement with an exponential likelihood) give a result very close to the full-information one.
Model Value ${\sigma^{+}}$ ${\sigma^{-}}$
Full information 1.1325 0.6225 0.3598
Poisson 1.1285 0.6282 0.3533
Cubic 1.1074 NA 0.3433
Quartic 1.1335 0.6243 0.3637
Logarithmic 1.1319 0.6237 0.3586
Linear sigma 1.1323 0.6213 0.3604
Linear variance 1.1318 0.6249 0.3577
PDG method 1.1318 0.6249 0.3613
Skew normal 1.1494 0.6146 0.3764
Wrong 1.1325 0.7261 0.3326
Table 9: Results from combining two lifetime experiments.
@@
These results are shown in Table 9 together with those from some other models. The Edgeworth model cannot cope with asymmetries this large, and the cubic does not find a positive error due to the turnover in the curve. It can be seen that the logarithmic and PDG models also do well, the quartic is less good, and the Poisson and skew normal models perform relatively poorly. The final row, labelled “wrong”, takes the mean of the two results and combines the positive and negative errors separately in quadrature, and it can be seen that its error estimates are seriously discrepant.
4.1.3 Combining Poisson counts: combination of results using likelihoods
@@
Suppose a counting experiment sees 5 events in an hour. The result is quoted, using $\Delta\ln L=-{1\over 2}$ errors even though this is a case where the full Neyman errors could be given, as ${5}^{+2.581}_{-1.916}$. This continues for another hour and, as it happens, 5 events are again seen. The total gives a result ${10}^{+3.504}_{-2.838}$ and with the knowledge we have of the way the experiment has been done, we can estimate the number of events per hour by dividing this by 2 to get ${5}^{+1.752}_{-1.419}$.
@@
${5}^{+2.581}_{-1.916}$ ${5}^{+2.581}_{-1.916}$ ${5.000}^{+1.737}_{-1.408}$ ${5.000}^{+1.748}_{-1.415}$ ${5.000}^{+1.732}_{-1.395}$ ${5.000}^{+1.719}_{-1.399}$
${6}^{+2.794}_{-2.128}$ ${4}^{+2.346}_{-1.682}$ ${4.998}^{+1.778}_{-1.432}$ ${5.000}^{+1.759}_{-1.425}$ ${5.009}^{+1.766}_{-1.460}$ ${4.992}^{+1.815}_{-1.440}$
${7}^{+2.989}_{-2.323}$ ${3}^{+2.080}_{-1.416}$ ${5.038}^{+1.937}_{-1.530}$ ${5.009}^{+1.794}_{-1.456}$ ${5.002}^{+1.744}_{-1.594}$ ${5.107}^{+2.147}_{-1.651}$
${8}^{+3.171}_{-2.505}$ ${2}^{+1.765}_{-1.102}$ ${5.401}^{+2.368}_{-1.826}$ ${5.054}^{+1.856}_{-1.516}$ ${4.689}^{+1.635}_{-1.589}$ ${5.282}^{+1.277}_{-1.707}$
${9}^{+3.342}_{-2.676}$ ${1}^{+1.358}_{-0.6983}$ ${7.348}^{+3.149}_{-2.549}$ ${5.201}^{+1.942}_{-1.605}$ ${3.633}^{+1.426}_{-1.414}$ ${3.209}^{+0.5942}_{-0.8277}$
Table 10: Combining results from two samples from the same Poisson distribution. The ideal result is ${5.000}^{+1.752}_{-1.419}$.